OpenStudy (anonymous):

An athlete runs with velocity 40 km/h for 5 minutes, 40 km/h for the next 2 minutes, and 12 km/h for another 3 minutes. Compute the total distance traveled.

OpenStudy (zzr0ck3r):

distance = rate * time you have three legs first leg: rate = 40 time = 5 distance = rate * time so distance for the first leg is 200 do this for the other two legs and add them together

OpenStudy (anonymous):

@zzr0ck3r that was incorrect when I tried it.:(

OpenStudy (zzr0ck3r):

what did you get for the second leg?

OpenStudy (anonymous):

40*2=80 then 12*3=36 added all together 316

OpenStudy (zzr0ck3r):

o sorry you need to readjust for the rate because its by the hour so 5 min = 5/60 of an hour so leg 1 40(5/60)

OpenStudy (anonymous):

I forgot to divide I got it thanks!